OpenStudy (anonymous):

solve please!!! |x + 4| = 4

OpenStudy (anonymous):

x=0

OpenStudy (anonymous):

|x + 4| = 4 x + 4 = 4 -4 -4 x = 0 |x + 4| = 4 x + 4 = -4 - 4 -4 x = -8